# Quadratic Modelings of Syndrome Decoding

### v1

@@ 5 Complexity Analysis of Modeling 2
Since Theorem 5.2 considers the set $Q^{\mathrm{top}}\cup P^{\mathrm{top}}=S^{\mathrm{top}}\setminus L_{H}^{\mathrm{top}}$, it only gives an upper bound to the degree of regularity $d_{\mathrm{reg}}(S)$.
@@ 5.1 Experimental results
We performed several experiments for Modeling 2 taking as input both random and Goppa codes, and we obtained a solving degree which is much smaller than the upper bound for the degree of regularity computed in Theorem 5.2. This results in a much lower complexity estimate. To give a taste of our results, we report a selection of our experiments in Table 3. The MAGMA code used for our experiments can be found at this link.
$n$ $k$ $t$ Code Type # Lin Eqs # Quad Eqs # Vars $d_{\mathrm{reg}}$ SR $d_{\mathrm{reg}}$ SD
8 2 2 Goppa 17 38 24 $\leq 8$ 3 2
10 5 4 Random* 20 67 40 $\leq 14$ 5 3
16 8 2 Goppa 27 78 48 $\leq 16$ 5 3
20 10 5 Random* 35 137 80 $\leq 29$ 7 3
30 15 7 Random* 50 207 120 $\leq 44$ 10 4
32 12 4 Goppa 57 221 128 $\leq 47$ 10 3
32 17 3 Goppa 50 158 96 $\leq 32$ 5 4
32 22 2 Goppa 45 158 96 $\leq 32$ 7 3
40 20 8 Random* 67 356 160 $\leq 78$ 16 5
50 30 5 Random 75 347 200 $\leq 74$ 15 4
50 40 4 Random 65 347 200 $\leq 74$ 17 4
64 52 2 Goppa 79 318 192 $\leq 64$ 14 4
64 40 4 Goppa 93 445 256 $\leq 95$ 19 4
64 16 8 Goppa 119 572 320 $\leq 126$ 21 4
@@
Table 3: This table gives information from experiments of $\mathbb{F}_{2}$-linear codes using Modeling 2 for the ESDP. The values in the SD column represent the highest step degree achieved when directly computing the Gröbner basis of the system in MAGMA. This is typically regarded as a proxy for the solving degree $d_{\mathrm{sol}}$. The SR $d_{reg}$ column gives the degree of regularity of a semi-regular system of equations with the associated number of linear equations, quadratic equations, and variables, using [3, Corollary 3.3.8]. The values in $d_{\mathrm{reg}}$ column are upper bounds for the degree of regularity of the system as provided by Theorem 5.2 and Remark 7. Random codes with “*” are decoding challenges from https://decodingchallenge.org/syndrome, with a number of errors slightly above Gilbert-Varshamov distance. Several solutions are indeed found. The other random code instances are below GV distance instead. Instances with “Goppa” Code Type are random full-length binary Goppa codes with a number of errors equal to the Goppa polynomial degree. These parameters have also been tested with random codes and always provide the same solving degree in the two cases.
@@ 6.1 Construction of the Equations
A difficulty arising from a direct generalization to large fields of the previous approach is the update of the weight registers, i.e. of the vectors $\mathbf{y}_{i}$’s. In order to overcome this limitation, we introduce a different strategy for encoding the partial weights. More precisely, we substitute their binary expansion with the representation of finite field elements through companion matrices. This approach associates intervals of integers with powers of matrices and naturally allows for the choice of different trade-offs between the number of variables and finite field size.
@@
We also recall that the order of a nonzero polynomial $f$ with $f_{0}=0$ is the least positive integer $e$ such that $f(x)\mid x^{e}-1$. A polynomial in $\mathbb{F}_{q}[x]$ of degree $d$ is said primitive if it is monic, $f(0)\neq 0$, and $\mathsf{ord}(f)=q^{d}-1$. Such polynomials can be found from the factorization of $x^{q^{d}-1}-1$.
The theory of linear recurring sequences says that the least period of the sequence of vectors $\mathbf{y}_{0},\mathbf{C}_{f}\mathbf{y}_{0},\mathbf{C}_{f}^{2}\mathbf{y}_{0},\dots$, for some nonzero $\mathbf{y}_{0}$ and companion matrix with $f_{0}\neq 0$, is periodic with least period equal to the order of $f$, when the latter is irreducible (cf. [28, Theorem 6.28]). Therefore, by choosing $f$ primitive, we obtain a sequence of vectors $\mathbf{y}_{0},\mathbf{C}_{f}\mathbf{y}_{0},\mathbf{C}_{f}^{2}\mathbf{y}_{0},\dots$ of maximal order $q^{d}-1$. On the other hand, the choice of $\mathbf{y}_{0}$ does not seem to affect any property of our modeling. Without loss of generality, we thus fix the initial state vector
$\mathbf{y}_{0}=\begin{pmatrix}1\\ 0\\ \vdots\\ 0\end{pmatrix}$ (14)
from now on.
As already anticipated, our strategy allows different ways to choose the number of variables used for the Hamming weight computation encoding, at the cost of working over more or less large field extensions. More precisely, take
@@
and let $r_{1},r_{2}$ be two positive integers such that $m\leq r_{1}r_{2}$.
Then, let $f\in\mathbb{F}_{q^{r_{1}}}[x]$ be a primitive polynomial of degree $r_{2}$ and $\mathbf{C}_{f}\in\mathbb{F}_{q^{r_{1}}}^{r_{2}\times r_{2}}$ its companion matrix. For convenience sake, we will use the column vector notation for the $Y_{j}$’s blocks of variables, i.e.
$Y_{j}=\begin{pmatrix}y_{j,1}\\ \vdots\\ y_{j,r_{2}}\end{pmatrix}.$
@@
    $Y_{1}-(1-z_{1})\cdot\mathbf{y}_{0}-z_{j}\cdot\mathbf{C}_{f}\cdot\mathbf{y}_{0}$ (15)
@@
Each element in the set corresponds to $r_{2}$ polynomial equations over $\mathbb{F}_{r_{1}}$, as each entry of the length-$r_{2}$ vectors must be equal to 0.
The next proposition shows that the Hamming weight of $\mathbf{z}$ is correctly computed.
@@
$\mathbf{y}_{j}=C_{f}^{z_{j}}\cdot\mathbf{y}_{j-1}=\mathbf{C}_{f}^{z_{j}}\mathbf{C}_{f}^{\mathsf{wt}(\pi_{j-1}(\mathbf{z}))}\mathbf{y}_{0}=\mathbf{C}_{f}^{\mathsf{wt}(z_{j})+\mathsf{wt}(\pi_{j-1}(\mathbf{z}))}\mathbf{y}_{0}=\mathbf{C}_{f}^{\mathsf{wt}(\pi_{j}(\mathbf{z}))}\mathbf{y}_{0}.$
@@
Substituting $\mathbf{z}$ in (15) uniquely determines all the values $y_{1,1},\dots,y_{1,r_{2}}$ by linear equations of the form $y_{1,i}=c_{i}$. Moreover, substituting $y_{j-1,1},\dots,y_{j-1,r_{2}}$ in (16) recursively determines all the values $y_{j,1},\dots,y_{j,r_{2}}$ in a similar manner. Therefore, if $\bar{\mathbf{z}}=\tilde{\mathbf{z}}$ are the projections over the last $n$ coordinate of two solutions, then also $(\bar{\mathbf{y}},\bar{\mathbf{z}})=(\tilde{\mathbf{y}},\tilde{\mathbf{z}})$, which concludes the proof. ∎
@@ 6.3 The Dimension of the Variety Associated with Modeling 5
A special case is given by the codeword finding problem, i.e. where the syndrome $\mathbf{s}$ is the zero vector. Here the the null vector is a solution of Problem 1 and leads to $\binom{n}{t}$ solutions, thus likely increasing a lot the solving degree and the cost of the Gröbner basis computation. We can get rid of all these solutions by fixing one variable $z_{i}=1$, thus forcing any solution to have weight at least 1 and removing the null vector. If the target solution has weight $t$, then the guess has success with probability $t/n$. We will discuss this strategy in more detail at the end of this subsection. ∎
@@
Let $\mathbb{K}$ be an algebraically closed field (we will apply the following definitions and results to $\mathbb{K}=\overline{\mathbb{F}_{q}}$). An affine variety $V$ is the zero locus in $\mathbb{K}^{m}$ of a proper ideal $I$ of the polynomial ring $\mathbb{K}[x_{1},\dots,x_{m}]$. We say that $V$ is irreducible if it is not possible to write $V=V_{1}\cup V_{2}$ where $V_{1},V_{2}\subsetneq V$ are two proper subvarieties. Irreducibility of $V$ is equivalent to the corresponding ideal $I$ being prime. The Krull dimension or simply the dimension of a variety $V$ is defined as the maximal length $d$ of the chains $V_{0}\subsetneq V_{1}\subsetneq\cdots\subsetneq V_{d}$, of distinct nonempty irreducible subvarieties of $V$. This is also equivalent to the supremum of the lengths of all chains of prime ideals containing the defining ideal $I$ of $V$. For example, the Krull dimension of an affine linear space $\mathcal{L}$ generated by $a$ linearly independent affine linear polynomials $L_{1},\dots,L_{a}$ is precisely $m-a$. This can be seen by completing the polynomials to a maximal linearly independent system of $m$ equations (in $m$ variables) $L_{1},\dots,L_{a},L_{a+1},\dots,L_{m}$ and then considering the following maximal chain of prime ideals
@@
Let $\mathcal{C}$ be an $[n,k]$ code with parity-check matrix $\mathbf{H}\in\mathbb{F}_{q}^{(n-k)\times n}$. Then the set of solutions of the Problem 1 with target weight $t$ and syndrome $\mathbf{s}$ for the code $\mathcal{C}$ is the finite union of irreducible components, namely
@@
On the other hand, for any solution $\mathbf{x}\in\mathbb{F}_{q}^{n}$ to the decoding problem, let $S$ be a set of cardinality $t$ containing the support of $\mathbf{x}$. Then, $\mathbf{x}\in\{\mathsf{pad}_{S}(\mathbf{x})\in\mathbb{F}_{q}^{n}\mid\mathbf{H}_{S}\mathbf{x}=\mathbf{s}\}$. Finally, all the sets $\{\mathsf{pad}_{S}(\mathbf{x})\in\mathbb{F}_{q}^{n}\mid\mathbf{H}_{S}\mathbf{x}=\mathbf{s}\}$ are irreducible being linear subspaces. ∎
@@
Let us fix the support $S$ of $t$ possible error positions. By Remark 10, the Krull dimension of the irreducible components coincides with their dimensions as vector spaces. The case study of the set of solutions of
@@
  • If $t>n-k$:
    • if $\mathsf{rk}(\mathbf{H}_{S})=n-k\Rightarrow$ the variety has dimension $t-n+k$;
    • if $\mathsf{rk}((\mathbf{H}_{S}\mid\mathbf{s}))=\mathsf{rk}(\mathbf{H}_{S})<n-k\Rightarrow$ the variety has dimension $t-\mathsf{rk}(\mathbf{H}_{S})$;
    • if $\mathsf{rk}((\mathbf{H}_{S}\mid\mathbf{s}))>\mathsf{rk}(\mathbf{H}_{S})<n-k\Rightarrow$ the variety is empty;
  • If $t=n-k$:
    • if $\mathsf{rk}((\mathbf{H}_{S}\mid\mathbf{s}))=\mathsf{rk}(\mathbf{H}_{S})=t\Rightarrow$ the variety has a unique element, and it belongs to $\mathbb{F}_{q}^{n}$;
    • if $\mathsf{rk}((\mathbf{H}_{S}\mid\mathbf{s}))>\mathsf{rk}(\mathbf{H}_{S})=t\Rightarrow$ the variety is empty;
    • if $\mathsf{rk}((\mathbf{H}_{S}\mid\mathbf{s}))=\mathsf{rk}(\mathbf{H}_{S})<t\Rightarrow$ the variety has dimension $t-\mathsf{rk}(\mathbf{H}_{S})$;
    • if $\mathsf{rk}((\mathbf{H}_{S}\mid\mathbf{s}))>\mathsf{rk}(\mathbf{H}_{S})<t\Rightarrow$ the variety is empty;
  • If $t<n-k$:
    • if $\mathsf{rk}(\mathbf{H}_{S})=t\Rightarrow$ the variety has a unique element, and it belongs to $\mathbb{F}_{q}$;
    • if $\mathsf{rk}((\mathbf{H}_{S}\mid\mathbf{s}))=\mathsf{rk}(H_{S})<t\Rightarrow$ the variety has dimension $t-\mathsf{rk}(H_{S})$;
    • if $\mathsf{rk}((\mathbf{H}_{S}\mid\mathbf{s}))>\mathsf{rk}(H_{S})<t\Rightarrow$ the variety is empty.
We observe that the solution vector $\mathbf{x}$ can have 0 entries, i.e. a weight smaller than $t$. This will increase the solutions of the algebraic system but not the dimension of the variety. Therefore, by Remark 10, the dimension of the solutions set is obtained as the maximum dimension over all the irreducible components corresponding to some $S$ for which $\mathsf{rk}((\mathbf{H}_{S}\mid\mathbf{s}))=\mathsf{rk}(\mathbf{H}_{S})$:
@@
Let us now consider the case of a zero-dimensional ideal. Then for any choice of $S$, there is at most a solution and it must belong to $\mathbb{F}_{q}^{n}$. Hence, all solutions belong to $\mathbb{F}_{q}^{n}$. ∎
@@
It readily follows from Propositions 10 and 5 and the fact that each solution of Problem 1 corresponds to a finite number of solutions of Modeling 5. ∎

### v2

@@ 5 Complexity Analysis of Modeling 2
Since Theorem 5.2 considers the set $Q^{\mathrm{top}}\cup P^{\mathrm{top}}=S^{\mathrm{top}}\setminus L_{\mathbf{H}}^{\mathrm{top}}$, it only gives an upper bound to the degree of regularity of $S$, that is
$d_{\mathrm{reg}}(S)\leq\left\lceil\frac{n-1}{2}\right\rceil+(\ell_{t}-1)\left\lceil\frac{n-2}{2}\right\rceil+1.$
@@ 5.1 Experimental results
We performed several experiments for Modeling 2 taking as input both random and Goppa codes, and we obtained a solving degree which is much smaller than the upper bound for the degree of regularity computed in Theorem 5.2. This results in a much lower complexity estimate. We provide a selection of our experiments in Table 3. The MAGMA code used for our experiments can be found here.
$n$ $k$ $t$ Code Type # Eqs # Vars $d_{\mathrm{reg}}$ SR $d_{\mathrm{reg}}$ $d_{\mathrm{M}}$ Prange Modeling 2
8 2 2 Goppa (17,38) 24 $\leq 8$ 3 2 9 23
10 5 4 Random* (20,67) 40 $\leq 14$ 5 3 10 38
16 8 2 Goppa (27,78) 48 $\leq 16$ 5 3 12 40
20 10 5 Random* (35,137) 80 $\leq 29$ 7 3 12 46
30 15 7 Random* (50,207) 120 $\leq 44$ 10 4 15 65
32 12 4 Goppa (57,221) 128 $\leq 47$ 10 3 16 52
32 17 3 Goppa (50,158) 96 $\leq 32$ 5 4 16 61
32 22 2 Goppa (45,158) 96 $\leq 32$ 7 3 15 48
40 20 8 Random* (67,356) 160 $\leq 78$ 16 5 17 88
50 30 5 Random (75,347) 200 $\leq 74$ 15 4 20 73
50 40 4 Random (65,347) 200 $\leq 74$ 17 4 14 73
64 52 2 Goppa (79,318) 192 $\leq 64$ 14 4 18 72
64 40 4 Goppa (93,445) 256 $\leq 95$ 19 4 21 77
64 16 8 Goppa (119,572) 320 $\leq 126$ 21 4 19 81
@@
Table 3: This table shows experimental results for $\mathbb{F}_{2}$-linear codes using Modeling 2 for the ESDP. The number of equations is given in the format (#linear equations, #quadratic equations). The values in the $d_{\mathrm{M}}$ column represent the smallest degree D such that MAGMA function GroebnerBasis(F,D) gives the Gröbner basis, i.e. highest step degree achieved when directly computing the Gröbner basis of the system in MAGMA, but ignoring the unnecessary steps in high degree that MAGMA F4 algorithm may do to insure termination. The SR $d_{reg}$ column gives the degree of regularity of a semi-regular system of equations with the associated number of linear equations, quadratic equations, and variables, using [3, Corollary 3.3.8]. The values in $d_{\mathrm{reg}}$ column are upper bounds for the degree of regularity of the system as provided by Theorem 5.2 and Remark 8. Random codes with “*” are decoding challenges from https://decodingchallenge.org/syndrome, with a number of errors slightly above Gilbert-Varshamov distance. The other random code instances are below GV distance instead. Instances with “Goppa” Code Type are random full-length binary Goppa codes with a number of errors equal to the Goppa polynomial degree. The Prange and Modeling 2 columns state the $\log_{2}$ of the complexities of the Prange algorithm (Esser-Bellini estimator [24]) and Gröbner basis computations (1) with $d_{\mathrm{M}}$, respectively.
5.1.1 Comparison with combinatorial attacks.
We also compare the complexity of solving ESDP using Gröbner basis techniques with the more traditional combinatorial attack proposed by Prange. The latter is widely regarded as a reference algorithm for decoding linear codes in the Hamming metric, and forms the basis of many state-of-the-art attacks on code-based cryptosystems.
However, Gröbner basis computations are known to be computationally expensive, particularly as the number of variables and the degree of the polynomials grow. In contrast, combinatorial methods like Prange exploit structured randomization and search to solve the problem more efficiently, albeit with large memory requirements.
In Table 3, the complexity of the Prange algorithm has been computed using the Esser-Bellini cryptographic estimator [24], see also this link. On the other hand, the complexity of solving the polynomial system of Modeling 2 with Gröbner bases methods has been computed using formula (1) with $N=n(\lfloor\log_{2}(t)\rfloor+2)$, the highest step degree $d_{M}$ achieved in the MAGMA experiments as a proxy for the solving degree $d_{\mathrm{sol}}$ and the conservative matrix multiplication constant $\omega=2.807$.
These results confirm that direct combinatorial attacks like Prange outperform algebraic methods when solving the syndrome decoding problem, particularly for parameters of practical interest in code-based cryptography. Despite this, algebraic approaches remain valuable from a theoretical perspective and may offer insights into alternative solution strategies that could be leveraged in other contexts. We include this comparison to emphasize that the purpose of our Gröbner basis approach is not to compete with combinatorial attacks in efficiency, but rather to provide an alternative algebraic perspective on the syndrome decoding problem. Such perspectives can contribute to understanding the hardness of related problems in post-quantum cryptography.
@@ 6.1 Construction of the Equations
A difficulty arising from a direct generalization to large fields of the previous approach is the update of the weight registers, i.e. of the vectors $\mathbf{y}_{i}$’s. In order to overcome this limitation, we introduce a different strategy for encoding the partial weights. More precisely, we substitute their binary expansion with vectors from a linear recurring sequence over an extension of $\mathbb{F}_{q}$. As we will see, this approach naturally allows for the choice of different trade-offs between the number of variables and finite field size.
@@
We also recall that the order of a non-constant polynomial $f$ with $f_{0}\neq 0$ is the least positive integer $e$ such that $f(x)\mid x^{e}-1$. A polynomial in $\mathbb{F}_{q}[x]$ of degree $d$ is said primitive if it is monic, $f(0)\neq 0$, and $\mathsf{ord}(f)=q^{d}-1$. Such polynomials can be found from the factorization of $x^{q^{d}-1}-1$.
The theory of linear recurring sequences says that the sequence of vectors $\mathbf{y}_{0},\mathbf{C}_{f}\mathbf{y}_{0},\mathbf{C}_{f}^{2}\mathbf{y}_{0},\dots$, for some nonzero $\mathbf{y}_{0}$ and companion matrix with $f_{0}\neq 0$, is periodic with least period equal to the order of $f$, when the latter is irreducible (cf. [29, Theorem 6.28]). Therefore, by choosing $f$ primitive, we obtain a sequence of vectors $\mathbf{y}_{0},\mathbf{C}_{f}\mathbf{y}_{0},\mathbf{C}_{f}^{2}\mathbf{y}_{0},\dots$ of maximal order $q^{d}-1$. On the other hand, the choice of $\mathbf{y}_{0}$ does not seem to affect any property of our modeling. Without loss of generality, from now on we fix the initial state vector
$\mathbf{y}_{0}=\begin{pmatrix}1&0&\cdots&0\end{pmatrix}^{\top}.$ (14)
By tuning the values $r_{1}$ and $r_{2}$ we can choose the number of variables used for the Hamming weight computation encoding, at the cost of working over more or less large field extensions. More precisely, take
@@
and let $r_{1},r_{2}$ be two positive integers such that $m\leq r_{1}r_{2}$. Then, let $f\in\mathbb{F}_{q^{r_{1}}}[x]$ be a primitive polynomial of degree $r_{2}$ and $\mathbf{C}_{f}\in\mathbb{F}_{q^{r_{1}}}^{r_{2}\times r_{2}}$ its companion matrix. For convenience sake, we will use the column vector notation for the $Y_{j}$’s blocks of variables, i.e.
$Y_{j}=\begin{pmatrix}y_{j,1}&\cdots&y_{j,r_{2}}\end{pmatrix}^{\top}.$
@@
    $Y_{1}-(1-z_{1})\cdot\mathbf{y}_{0}-z_{1}\cdot\mathbf{C}_{f}\cdot\mathbf{y}_{0}$ (15)
@@
Using this approach, $r_{1}$ determines the finite field over which we define the resulting polynomial system (and thus the Multivariate Quadratic Problem instance), while $r_{2}$ determines the number of variables required for the weight computation encoding (which is strictly linked to the computational complexity). Observe that, together, equations (15) and (16) correspond to $nr_{2}$ polynomial equations over $\mathbb{F}_{q^{r_{1}}}$ in $nr_{2}+n$ variables. In several cases, working with a small amount of variables (namely, $r_{1}$ large and $r_{2}$ small) is preferable, but there are some instances in which working over small finite fields with a large number of variables (i.e. $r_{1}$ small and $r_{2}$ large) is advantageous.
The next result shows that the Hamming weight of $\mathbf{z}$ is correctly computed.
@@
$\mathbf{y}_{j}=\mathbf{C}_{f}^{z_{j}}\cdot\mathbf{y}_{j-1}=\mathbf{C}_{f}^{z_{j}}\mathbf{C}_{f}^{\mathsf{wt}(\pi_{j-1}(\mathbf{z}))}\mathbf{y}_{0}=\mathbf{C}_{f}^{\mathsf{wt}(z_{j})+\mathsf{wt}(\pi_{j-1}(\mathbf{z}))}\mathbf{y}_{0}=\mathbf{C}_{f}^{\mathsf{wt}(\pi_{j}(\mathbf{z}))}\mathbf{y}_{0}.$
@@
Substituting $\mathbf{z}$ in (15) uniquely determines all the values $y_{1,1},\dots,y_{1,r_{2}}$ by linear equations of the form $y_{1,i}=c_{i}$. Moreover, substituting $y_{j-1,1},\dots,y_{j-1,r_{2}}$ in (16) recursively determines all the values $y_{j,1},\dots,y_{j,r_{2}}$ in a similar manner. Therefore, if $\bar{\mathbf{z}}=\tilde{\mathbf{z}}$ are the projections over the last $n$ coordinates of two solutions, then also $(\bar{\mathbf{y}},\bar{\mathbf{z}})=(\tilde{\mathbf{y}},\tilde{\mathbf{z}})$, which concludes the proof. ∎
@@ 6.3 The Dimension of the Variety Associated with Modeling 5
A special case is given by the codeword finding problem, i.e. where the syndrome $\mathbf{s}$ is the zero vector. Here the null vector is a solution of Problem 1 and leads to $\binom{n}{t}$ solutions, thus likely increasing a lot the solving degree and the cost of a Gröbner basis computation. We can get rid of all these solutions by fixing one variable $z_{i}=1$, thus forcing any solution to have weight at least 1 and removing the null vector. If the target solution has weight $t$, then the guess has success with probability $t/n$. We will discuss this strategy in more detail at the end of this subsection.
@@
Let $\mathbb{K}$ be an algebraically closed field (we will apply the following definitions and results to $\mathbb{K}=\overline{\mathbb{F}_{q}}$). An affine variety $V$ is the zero locus in $\mathbb{K}^{m}$ of a proper ideal $I$ of the polynomial ring $\mathbb{K}[x_{1},\dots,x_{m}]$. We say that $V$ is irreducible if it is not possible to write $V=V_{1}\cup V_{2}$ where $V_{1},V_{2}\subsetneq V$ are two proper subvarieties. Irreducibility of $V$ is equivalent to the corresponding ideal $I$ being prime. The Krull dimension or simply the dimension of a variety $V$ is defined as the maximal length $d$ of the chains $V_{0}\subsetneq V_{1}\subsetneq\cdots\subsetneq V_{d}$, of distinct nonempty irreducible subvarieties of $V$. This is also equivalent to the supremum of the lengths of all chains of prime ideals containing the defining ideal $I$ of $V$. For example, the Krull dimension of an affine linear space $\mathcal{L}$ generated by $a$ linearly independent affine linear polynomials $L_{1},\dots,L_{a}$ is precisely $m-a$, that is its dimension as an affine space. This can be seen by completing the polynomials to a maximal linearly independent system of $m$ equations (in $m$ variables) $L_{1},\dots,L_{a},L_{a+1},\dots,L_{m}$ and then considering the following maximal chain of prime ideals
@@
Let $\mathcal{C}$ be an $[n,k]$ code with parity-check matrix $\mathbf{H}\in\mathbb{F}_{q}^{(n-k)\times n}$. Then the set of solutions of Problem 1 with target weight $t$ and syndrome $\mathbf{s}$ for the code $\mathcal{C}$ is the finite union of irreducible components, namely
@@
On the other hand, for any solution $\mathbf{x}\in\mathbb{F}_{q}^{n}$ to the decoding problem, let $S$ be a set of cardinality $t$ containing the support of $\mathbf{x}$. Then, $\mathbf{x}\in\{\mathsf{pad}_{S}(\mathbf{x})\in\mathbb{F}_{q}^{n}\mid\mathbf{H}_{S}\mathbf{x}=\mathbf{s}\}$. Finally, all the sets $\{\mathsf{pad}_{S}(\mathbf{x})\in\mathbb{F}_{q}^{n}\mid\mathbf{H}_{S}\mathbf{x}=\mathbf{s}\}$ are irreducible being affine linear spaces. ∎
@@
Let us fix the support $S$ of $t$ possible error positions. By Remark 11, the Krull dimension of the irreducible components coincides with their dimensions as affine linear spaces. The case study of the set of solutions of
@@
  • if $\mathsf{rk}((\mathbf{H}_{S}\mid\mathbf{s}))>\mathsf{rk}(\mathbf{H}_{S})\Rightarrow$ the variety is empty;
  • if $\mathsf{rk}((\mathbf{H}_{S}\mid\mathbf{s}))=\mathsf{rk}(\mathbf{H}_{S})\Rightarrow$ the variety has dimension $t-\mathsf{rk}(\mathbf{H}_{S})$. In particular, if the variety has dimension 0, i.e. $\mathsf{rk}(\mathbf{H}_{S})=t$, then it has a unique element, which belongs to $\mathbb{F}_{q}$.
By Remark 11, the dimension of the solutions set is obtained as the maximum dimension over all the irreducible components corresponding to some $S$ for which $\mathsf{rk}((\mathbf{H}_{S}\mid\mathbf{s}))=\mathsf{rk}(\mathbf{H}_{S})$:
@@
Let us now consider the case of a zero-dimensional variety. Then for any choice of $S$, there is at most a solution and it must belong to $\mathbb{F}_{q}^{n}$. Hence, all solutions belong to $\mathbb{F}_{q}^{n}$. ∎
@@
It readily follows from Propositions 11 and 5 and the fact that each solution of Problem 1 corresponds to a finite number of solutions of Modeling 5, thus it does not increase the dimension of the variety. ∎